# Towards Dataset Copyright Evasion Attack against Personalized Text-to-Image Diffusion Models

### v1

@@ VI Potential Limitations and Future Directions
Firstly, although CEAT2I does not require any additional fine-tuning beyond the standard model fine-tuning on watermarked datasets, it introduces extra computational overhead during the watermark removal process. Specifically, it relies on extracting intermediate feature representations to detect watermarked samples and identify triggers, which adds additional time and resource consumption. Moreover, although the watermark mitigation step is based on a closed-form concept erasure solution, which avoids model fine-tuning, it still requires extra computational effort to complete. A promising direction for future research is to further simplify the CEAT2I pipeline. The goal of future work could be to develop an end-to-end framework that automatically integrates detection, identification, and mitigation into a single lightweight process. Reducing the number of stages could significantly lower the computational burden, making the attack more practical and efficient for real-world red-teaming scenarios.
Secondly, our CEAT2I is designed specifically for T2I diffusion models, such as Stable Diffusion, which rely on the alignment between textual prompts and visual content. While these models currently dominate the generative image synthesis landscape, the broader generative AI ecosystem is rapidly evolving to include other modalities, such as text-to-video, text-to-3D, and text-image-language foundation models. In these settings, the architecture and modality differ significantly. The effectiveness of CEAT2I has not been validated outside the scope of image generation tasks. Adapting or generalizing CEAT2I to handle these models would require entirely new paradigms for watermarked sample detection, trigger localization, and watermark mitigation, possibly leveraging multimodal feature disentanglement. As such, a key direction for future work is to explore whether the foundational ideas behind CEAT2I, such as early convergence analysis and concept erasure, can be extended for other multimodal generative models. Doing so could help assess the robustness of ownership verification systems across diverse generative AI technologies and provide more holistic protection mechanisms for diverse data modalities.
@@ VII Conclusion
In this paper, we presented CEAT2I, a novel and effective copyright evasion attack targeting dataset ownership verification (DOV) in T2I diffusion models. While DOV techniques offered a promising solution for protecting datasets via backdoor-based watermarking, we demonstrated that they remain vulnerable to well-crafted evasion attacks. Our method leveraged three key components, including watermarked sample detection via feature convergence analysis, trigger identification through token-level ablation, and efficient watermark removal via closed-form model editing, which could neutralize both global and local watermarks without requiring additional fine-tuning. Extensive experiments across four DOV methods and three datasets showed that our CEAT2I significantly outperformed prior potential attack methods, effectively removing watermarks while preserving model fidelity and visual quality. Our findings revealed the pressing need to revisit assumptions about the robustness of DOV systems, and we hope that our CEAT2I will serve as a useful tool for stress-testing future ownership verification techniques in generative models.
Ethics Statement
This work aims to investigate the security vulnerabilities of DOV methods based on backdoor techniques in T2I diffusion models. All experiments with our proposed CEAT2I are conducted strictly within controlled laboratory environments, using only publicly available open-source datasets. We explicitly emphasize that CEAT2I is designed solely for research purposes to highlight potential risks and limitations in existing DOV mechanisms. We do not support the deployment of CEAT2I in real-world applications for malicious purposes. The ultimate goal of this work is to raise awareness among the community about the potential threats to DOV security and to call for more robust, reliable DOV methods in future designs.

### v2

@@ VI Potential Limitations and Future Directions
Firstly, although CEAT2I does not require any additional fine-tuning beyond the standard model fine-tuning on watermarked datasets, it introduces extra computational overhead during the watermark removal process. Specifically, it relies on extracting intermediate feature representations to detect watermarked samples and identify triggers, which adds additional time and resource consumption. A promising direction for future research is to further simplify the CEAT2I pipeline. The goal of future work could be to develop an end-to-end framework that automatically integrates detection, identification, and mitigation into a single lightweight process.
Secondly, our CEAT2I is designed specifically for T2I models, such as Stable Diffusion, which rely on the alignment between textual prompts and visual content. While these models currently dominate the generative image synthesis landscape, the broader generative AI ecosystem is rapidly evolving to include other modalities, such as text-to-video, text-to-3D, and text-image-language foundation models. In these settings, the architecture and modality differ significantly. The effectiveness of CEAT2I has not been validated outside the image generation tasks. As such, a key direction for future work is to explore whether the foundational ideas behind CEAT2I, such as early convergence analysis and concept erasure, can be extended for other multimodal generative models.
@@ VII Conclusion
In this paper, we presented CEAT2I, a novel and effective copyright evasion attack targeting DOV in T2I models. While DOV techniques offered a promising solution for protecting datasets via backdoor-based watermarking, we demonstrated that they remain vulnerable to well-crafted evasion attacks. Our method leveraged three key components, including watermarked sample detection via feature convergence analysis, trigger identification through token-level ablation, and efficient watermark removal via closed-form model editing. Extensive experiments across four DOV methods and three datasets showed that our CEAT2I significantly outperformed prior potential attack methods, effectively removing watermarks while preserving model fidelity and visual quality.
Ethics Statement. This work aims to investigate the security vulnerabilities of DOV methods based on backdoor techniques in T2I diffusion models. All experiments with our proposed CEAT2I are conducted strictly within controlled laboratory environments, using only publicly available open-source datasets. We emphasize that CEAT2I is designed solely for research purposes to highlight potential risks in existing DOV mechanisms. We do not support the deployment of CEAT2I in real-world applications for malicious purposes.
Acknowledgement
This work is supported in part by the National Natural Science Foundation of China under Grant 62571298 and Tsinghua SIGS KA Cooperation Fund.